# EigenLoRAx: Recycling Adapters to Find Principal Subspaces for Resource-Efficient Adaptation and Inference

### v3

@@ 3.2 Algorithm
$\lVert W_{\mathcal{E}}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}-\lVert{W^{*}}_{d+1}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}\leq\|W_{\mathcal{E}}-{W^{*}}_{d+1}\lVert^{2}_{F}$
@@ 4.4 Text-to-Image Image Generative Models
Failure Cases and Limitations Despite its advantages, EigenLoRAx has limitations. Figure 5 shows a failure case where the method fails to capture a key property of the desired image. While tasks may share a principal subspace, missing critical orthogonal components can degrade performance, especially if they were absent in the pretrained LoRAs used for extraction or if the chosen top $K$ components were suboptimal. In the latter case, empirical analysis of hyperparameters (Appendix B) can guide optimal $K$ selection. Additionally, our subspace augmentation method (Table 3) helps by iteratively sampling and orthogonalizing more components to recover missing subspace elements. A simple extension can further mitigate this issue by allowing a small number of rank-1 weights to be trainable outside the subspace. Another key limitation (Section 4.3) is the computational cost and instability of processing a large number of initial LoRAs. A continual learning approach building on our method could address this. Finally, our experiments did not explore layer-wise or weight matrix-level optimizations; we tested different $K$ values but kept them fixed across layers and for both A and B matrices. Additional failure cases are discussed in Section B.2.
@@ A Note on Prior Work
This paper works on some ideas initially introduced in the EigenLoRA work (Kaushik et al., 2025), which was rejected from ICLR’25 and remains available on OpenReview. Due to irreconcilable differences regarding research ethics and authorship integrity, the primary authors of this work chose to independently develop and extend their contributions. These contributions in the previous draft included the initial algorithm, all experimental results (excluding Section 4.2.2), as well as all ablation studies and additional experiments. The code used for these experiments was also developed solely by the first authors of this work and is publicly available. Since the prior work was neither accepted nor published, there are no concerns regarding self-plagiarism. We also acknowledge discussions and exchanges with some of the earlier authors, which, while not directly contributing to the technical developments of this work, provided context and background that shaped our understanding of the broader problem. We have taken extensive measures to ensure that this paper exclusively reflects the contributions of the authors listed, and any similarities beyond our own contributions are purely coincidental.

### v4

@@ 3.2 Algorithm
$\lVert W_{\mathcal{E}}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}-\lVert{W^{*}}_{d+1}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}$
$\leq\|W_{\mathcal{E}}-{W^{*}}_{d+1}\lVert^{2}_{F}$
@@
Why does the orthogonal principal component augmentation works?
Minimum Hyperspherical Energy
@@ 4.4 Text-to-Image Image Generative Models
Failure Cases and Limitations
Despite its advantages, EigenLoRAx has limitations. Figure 5 shows a failure case where the method fails to capture a key property of the desired image. While tasks may share a principal subspace, missing critical orthogonal components can degrade performance, especially if they were absent in the pretrained LoRAs used for extraction or if the chosen top $K$ components were suboptimal. In the latter case, empirical analysis of hyperparameters (Appendix B) can guide optimal $K$ selection. Additionally, our subspace augmentation method (Table 3) helps by iteratively sampling and orthogonalizing more components to recover missing subspace elements. A simple extension can further mitigate this issue by allowing a small number of rank-1 weights to be trainable outside the subspace. Another key limitation (Section 4.3) is the computational cost and instability of processing a large number of initial LoRAs. A continual learning approach building on our method could address this. Finally, our experiments did not explore layer-wise or weight matrix-level optimizations; we tested different $K$ values but kept them fixed across layers and for both A and B matrices. Additional failure cases are discussed in Section B.2.
@@ A Note on Prior Work
This paper reintroduces some ideas initially introduced by the first authors (of this work) in an unpublished work (Kaushik et al., 2025). Due to serious issues concerning research ethics, authorship integrity, and academic misconduct, the primary authors of this work (including the first author of both works) chose to independently develop and extend their contributions. These contributions in the previous draft included the original idea and initial algorithm, all experimental results (excluding Section 4.2.2), as well as all ablation studies and additional experiments. The code used for these experiments was also developed solely by the first authors of this work and is publicly available. As the prior work was not accepted or published, self-plagiarism is not a concern. We also acknowledge high-level discussions and exchanges with some of the earlier authors, who, while not directly contributing to the technical developments of this work, provided a limited context that shaped our understanding of the broader problem. We have taken extensive measures to ensure that this paper exclusively reflects the contributions of the authors listed, and any similarities beyond our own contributions are purely coincidental.